OpenStudy (anonymous):

\[\int\limits_{}^{} \frac{ x ^{2} }{ 6x ^{3} +9}\] what is the indefinite integral?

OpenStudy (anonymous):

u may use u-sub, let \(u=6x^3+9\) and see what happens

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[\int\frac{ x ^{2} }{ 6x ^{3} +9}dx\]\[=\frac{1}{3}\int\frac{1 }{ 6x ^{3} +9}d(x^3)\]\[=\frac{1}{3}\int\frac{1 }{ 6u +9}du\], u=x^3

OpenStudy (anonymous):

i don't see that as an answer :(

OpenStudy (anonymous):

You should try integrating it :)

OpenStudy (anonymous):

oh ok